OpenStudy (codysek98):

Find the slope of the curve at the indicated point. Use the limit of the difference quotient. y=7/(5+x) , x=8 I think the setup is ([7/(5+(8+h))]-[7/(5+8)])/h Once i get there and add up the denominators, I'm not sure how to go about finishing it.

OpenStudy (codysek98):

Okay, so I think I'm getting somewhere (maybe). I tried my best to continue and simplify and maybe get common denominators in my numerator. This is what I have; ([91/(169+13h)]-[(91+7h)/(169+13h)])/h

OpenStudy (codysek98):

I know the answer is: m=-7/169

OpenStudy (dumbcow):

your work looks correct so far. Just add the numerators together. Factor out an "h" This will cancel with "h" in denominator.

OpenStudy (holsteremission):

The difference quotient is \[\frac{y(x+h)-y(x)}{h}=\frac{\dfrac{7}{5+(x+h)}-\dfrac{7}{5+x}}{h}\]and so the difference quotient when \(x=8\) is \[\frac{y(x+h)-y(x)}{h}=\frac{\dfrac{7}{13+h}-\dfrac{7}{13}}{h}\]The slope is then \[\lim_{h\to0}\frac{\dfrac{7}{13+h}-\dfrac{7}{13}}{h}\]which you can compute by finding the LCD of the numerator and simplifying: \[\frac{\dfrac{7}{13+h}-\dfrac{7}{13}}{h}=\frac{\dfrac{7\times13}{13(13+h)}-\dfrac{7(13+h)}{13(13+h)}}{h}=\frac{-7h}{13h(13+h)}=-\frac{7}{13(13+h)}\]

OpenStudy (codysek98):

So I subtracted my numerators; 91-91+7h, so my numerator is 7h. I have (7h/(169+13h))/h How do I factor an h out of that equation?

OpenStudy (codysek98):

Also, how is it going to end up negative? The answer we're given is -7/169

OpenStudy (dumbcow):

use @HolsterEmission work as reference. Notice the bottom "h" can be put in denominator \[\rightarrow \frac{\frac{a}{b}}{c} = \frac{a}{bc}\] Then the h will cancel with h in term "-7h" in numerator

OpenStudy (dumbcow):

its negative, you forgot to distribute the negative ---> 91 - (91+7h) = 91 - 91 - 7h = -7h

OpenStudy (dumbcow):

The last step is applying the limit .... h =0

OpenStudy (codysek98):

Oh okay! I didn't realize you had to distribute the negative. I thought that since they had common denominators, you just combined them as in (91-91+7h). Thank you. So in @HolsterEmission's last step, you just plug 0 in for h?

OpenStudy (dumbcow):

yep

OpenStudy (codysek98):

Oh, awesome. Thank you!

OpenStudy (dumbcow):

yw fyi use parenthesis around multiple terms to help remind you to distribute

OpenStudy (codysek98):

Okay. Thank you!